zarkam21 (zarkam21):

Help with 7 part question

OpenStudy (rogerdodger6868):

where the question at?

zarkam21 (zarkam21):

@mathstudent55

OpenStudy (mathstudent55):

For part 1, use the hint, and replace P with the actual perimeter.

OpenStudy (mathstudent55):

Are you there?

zarkam21 (zarkam21):

Yes I am here now

OpenStudy (mathstudent55):

Do part 1 as I wrote above. What do you get?

zarkam21 (zarkam21):

2(l*W)

OpenStudy (mathstudent55):

That is not it. Do you see the hint in part 1? It gives you an equation for the perimeter of a rectangle. P = 2(L + W) Rewrite it as 2(L + W) = P Now replace P with the actual perimeter you are given.

zarkam21 (zarkam21):

27?

zarkam21 (zarkam21):

2(L+W)=27?

OpenStudy (mathstudent55):

Correct. That is Part 1.

OpenStudy (mathstudent55):

For Part II, take that equation and solve for W.

zarkam21 (zarkam21):

So take this same equation and subtract w?

zarkam21 (zarkam21):

Well nvm , i said that wrong

OpenStudy (mathstudent55):